In a sample that has 159 CFU's, with a 10 -9 dilution. What is the CFU's/ml if 0.1ml was plated in scientific notation?
Answer #1

CFU/ml = = Number of colonies / Volume plated    (in ml) *  Dilution factor .

CFU/ml = 159 / 0.1 * 10-9    = 159 * 1010   or    =   1.59 * 1012

* Therefore 1.59 * 1012 Colony forming Units are present in 1 ml of sample.